How many grams of water can be heated from 25.0oC to 36.7oC by the heat released from 78.0 g of copper that cools from 88.5oC to 32.8oC?

How many grams of copper can be cooled from 79.7 to 26.0 °C by the heat gained by 100.0 g of water that has an increase in temperature from 21.0 to 39.8 °C?

A vat of 5.25 kg of water underwent a decrease in temperature from 60.33 to 58.21 °C. How much energy in kilojoules left the water? Include the sign in your answer. (For this range of temperature, use a value of 4.18 J·g-1·°C-1 for the specific heat of water.)

In the reaction between formic acid (HCHO2) and sodium hydroxide, water and sodium formate (NaCHO2) are formed. To determine the heat of reaction, 75 mL of 1.07 M HCHO2 was placed in a coffee cup calorimeter at a temperature of 20.3 °C, and 45 mL of 1.78 M NaOH, also at 20.3 °C, was added. The mixture was stirred quickly with a thermometer, and its temperature rose to 24.1 °C. Write the balanced chemical equation for the reaction. Calculate the heat of reaction in joules. Assume that the specific heats of all solutions are 4.184 J g-1°C-1 and that all densities are 1.00 g mL-1. What is the heat of reaction per mole of acid (in units of kJ mol-1).

Write the balanced reaction that takes place:
Answer #1

1. How many grams of water can be heated from 25.0oC to 36.7oC by the heat released from 78.0 g of copper that cools from 88.5oC to 32.8oC?

Heat released by copper = mass of copper * specific heat capacity of copper * temperature change = 78.0 g * (0.385 J/ g 0C) * ( 32.8 - 88.5) 0 C = 1672.7 J

Heat gained by water = mass of water * specific heat capacity of water * temperature change = x g * (4.184 J/ g 0C) * ( 36.7 - 25) 0 C = 49x Joules

So, 49x = 1672.7

Or, x = 34.1

So, mass of water = 34.1 g

2. How many grams of copper can be cooled from 79.7 to 26.0 °C by the heat gained by 100.0 g of water that has an increase in temperature from 21.0 to 39.8 °C?

Heat gained by water = mass of water * specific heat capacity of water * temperature change = 100 g * (4.184 J/ g 0C) * ( 39.8 - 21) 0 C = 7866 Joules

Heat released by copper = mass of copper * specific heat capacity of copper * temperature change = x g * (0.385 J/ g 0C) * ( 26 - 79.7) 0 C = -20.7x Joules

So, 20.7x = 7866

Or, x = 380

So, mass of copper = 380 grams

3. A vat of 5.25 kg of water underwent a decrease in temperature from 60.33 to 58.21 °C. How much energy in kilojoules left the water? Include the sign in your answer. (For this range of temperature, use a value of 4.18 J·g-1·°C-1 for the specific heat of water.)

Mass of water = 5.25 kg = 5250 g

Heat released by water = mass of water * specific heat capacity of water * temperature change = 5250 g * (4.184 J/ g 0C) * ( 58.21 - 60.33) 0 C = - 46567 Joules = - 46.567 kJ

So, amount of heat released = 46.567 kJ

4. In the reaction between formic acid (HCHO2) and sodium hydroxide, water and sodium formate (NaCHO2) are formed. To determine the heat of reaction, 75 mL of 1.07 M HCHO2 was placed in a coffee cup calorimeter at a temperature of 20.3 °C, and 45 mL of 1.78 M NaOH, also at 20.3 °C, was added. The mixture was stirred quickly with a thermometer, and its temperature rose to 24.1 °C. Write the balanced chemical equation for the reaction. Calculate the heat of reaction in joules. Assume that the specific heats of all solutions are 4.184 J g-1°C-1 and that all densities are 1.00 g mL-1. What is the heat of reaction per mole of acid (in units of kJ mol-1).

Balanced equation: HCOOH (aq) + NaOH (aq) = HCOONa (aq) + H2O (l)

Mass of HCHO2 solution = volume * density = 75 mL * ( 1 g / mL) = 75 g

Mass of NaOH solution = volume * density = 45 mL * ( 1 g/ mL) = 45 g

Total mass of solution = (75 + 45) g = 120 g

Total heat gained by solution (Q) = mass of solution * specific heat capacity of solution * temperature change = 120 g * (4.184 J/ g 0C) * ( 24.1 - 20.3) 0 C = 1908 Joules

Moles of HCHO2 = volume in L * Molarity = (75/ 1000) L * 1.07 M = 0.08025 mole

Heat of reaction per mole of acid = 1908 J/ 0.08025 mole = 23776 J / mole = 23.776 kJ/ mole